OpenStudy (anonymous):

The graph below shows the fine that a library member pays based on the number of days a book is overdue: A graph is shown. The values on the x axis are 0, 2, 4, 6, 8. The values on the y axis are 0, 1.50, 3.00, 4.50, and 6.00. Points are shown on ordered pairs 0, 0 and 2, 1.50 and 4, 3.00 and 6, 4.50. These points are joined by a line. The label on the x axis is Days Overdue. The title on the y axis is Fine.Which statement best describes the point (0, 0) on the graph? No fine is paid if the book is returned by the due date. A fine of $0.75 is paid if the book returned by the due date.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

The number of pages that Ana, Hillary, Roger, and Juan can read in a day is shown below: Ana read 15% of her 46-page book. Hillary read 11% of her 72-page book. Roger read 12% of his 68-page book. Juan read 14% of his 69-page book. Who can read the greatest number of pages in a day? Ana Hillary Roger Juan

OpenStudy (michele_laino):

I think "No fine is paid if the book is returned by the due date"

OpenStudy (anonymous):

thx, could you help with the other one pls??

OpenStudy (anonymous):

???

OpenStudy (anonymous):

The number of pages that Ana, Hillary, Roger, and Juan can read in a day is shown below: Ana read 15% of her 46-page book. Hillary read 11% of her 72-page book. Roger read 12% of his 68-page book. Juan read 14% of his 69-page book. Who can read the greatest number of pages in a day? Ana Hillary Roger Juan

OpenStudy (michele_laino):

from your data, we have: Ana--->690/100 Hillary--->792/100 ROger--->816/100 Juan--->966/100 so: Juan, because 966>816>792>690

OpenStudy (anonymous):

thx ur awesome :D

OpenStudy (michele_laino):

thank you!:)

OpenStudy (anonymous):

Meg conducted a survey of four school cafeterias to find the number of students who like sandwiches for lunch. The results of her survey are recorded in the table below: School Cafeteria Survey School Total Number of Students in the Cafeteria Number of Students Who Liked Sandwiches A 38 12 B 48 9 C 26 10 D 24 8 Which school has the greatest percentage of students who like sandwiches for lunch? School A School B School C School D

OpenStudy (anonymous):

i hope u can understand the table @Michele_Laino

OpenStudy (anonymous):

srry

OpenStudy (michele_laino):

no problem, it is okay!

OpenStudy (michele_laino):

I think that your answer is: school A--->12/38=31.6% school B--->9/48=18.7% school C--->10/26=38.5% school D--->8/24=33.3% so: school C

OpenStudy (anonymous):

thx

OpenStudy (michele_laino):

thanks!!!!!

OpenStudy (anonymous):

Convert fraction with numerator 7 and denominator 8 to a decimal. 0.375 0.875 0.925 0.965

OpenStudy (michele_laino):

we have: \[\frac{ 7 }{ 8 }=\frac{ 70/8 }{ 10 }=\frac{ 8+6/8 }{ 10 }=\frac{ 8+3/4 }{ 10 }=\] \[=\frac{ 8 }{ 10 }+\frac{ 3 }{ 4 }*\frac{ 1 }{ 10 }=0.8+0.75*\frac{ 1 }{ 10 }=\] \[=0.8+0.075=0.875\] so the second option

OpenStudy (anonymous):

Which unit rate is equivalent to 14 miles per gallon? two gallons over thirty two miles thirty two miles over two gallons three miles over forty two gallons forty two gallons over three miles

OpenStudy (michele_laino):

I think forty two miles over three gallons

OpenStudy (michele_laino):

even if my answer not appears in your options

OpenStudy (anonymous):

thx :D

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

Which expression is equivalent to 2(m − 5) + 6? 2m − 1 2m + 22 2m − 4 2m + 16

OpenStudy (michele_laino):

we have: \[2(m-5)+6=2m-10+6=2m-4\] so: the third option

OpenStudy (anonymous):

thx :D what do you think about this one @Michele_Laino ??

OpenStudy (anonymous):

The temperature at a mountain base camp was −2 degrees Celsius on Thursday. Friday morning, the temperature was 1 degree Celsius lower than what it was on Thursday. By Friday evening, the temperature was 3 degrees Celsius lower than what it was that morning. What was the temperature at the base camp Friday evening? −9 degrees Celsius −8 degrees Celsius −7 degrees Celsius −6 degrees Celsius

OpenStudy (anonymous):

the ??? are -

OpenStudy (michele_laino):

we have: Thursday--->-2 Friday morning--->-2-1=-3 Friday evening--->-3-3=-6 so the fourth option

OpenStudy (anonymous):

thx

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

Where is the number 3 − 7 located on a horizontal number line? 4 units to the left of 3 7 units to the left of 3 4 units to the right of 3 7 units to the right of 3

OpenStudy (anonymous):

3-7

OpenStudy (anonymous):

after this question ill open another one :D

OpenStudy (michele_laino):

since 3-7=-4, then it is located at 7 units to the left of 3

OpenStudy (michele_laino):

so the second option

OpenStudy (michele_laino):